OpenStudy (anonymous):

Evaluate limit (x+1)/(2x-6)/(x^2-4) as x approaches 2

OpenStudy (nowhereman):

You have to place more parentheses, to give that term a meaning!

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

its 3 things dividing each other

OpenStudy (anonymous):

so on top is \[((x+1)\div(2x-6))\div(x^2-4)\]

OpenStudy (anonymous):

is that better?

OpenStudy (anonymous):

its an improper fraction.

OpenStudy (nowhereman):

Yes. The limit does not exist, there is a pole at x=2.

OpenStudy (anonymous):

Right, there should be a removable discontinuity because the answer is -1/8

OpenStudy (nowhereman):

No, \(x^2 - 4\) tends to 0 and as it is in the denominator, plus the numerator is continuous at x=2, there must be a pole.

OpenStudy (anonymous):

So, the answer does not exist? Weird, because the answer key says that its -1/8.

OpenStudy (nowhereman):

Then the formula must be different.

OpenStudy (anonymous):

I simplified it a bit.

OpenStudy (anonymous):

The original formula is...

OpenStudy (anonymous):

\[((x/2)+(1/(x-3))/(x^2-4)\]

OpenStudy (nowhereman):

Then you must have made a mistake, because here the numerator tends to 0 too.

OpenStudy (anonymous):

I combined the numerator to find a common denominator for them.

OpenStudy (anonymous):

That's the original expression.

OpenStudy (anonymous):

I'm assuming I have to manipulate it to cancel out x-2 or x+2.

OpenStudy (anonymous):

and remove the discontinuity

OpenStudy (nowhereman):

Yes, \(x-2\). But it should work: \[\frac x 2 + \frac 1 {x-3} = \frac{x(x-3) + 2}{2(x-3)} = \frac{x^2-3x+2}{2(x-3)} = \frac{(x-2)(x-1)}{2(x-3)}\]

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

nevermind.